A wheel starting from rest is uniformly accelerated at 4 rads/s² for 10 seconds, It is allowed to rotate uniformly for the next 10 seconds and is finally brought to rest in the next 10 seconds. Find the total angle rotated by the wheel?

Grade:12th pass

1 Answers

Khimraj
5 years ago
Time duration t = 10 sec.
angle rotation = θ
Here the rotation in the first and third part will be same.
Hence Total angle rotation,
{   }
=θ+θ'+θ      {here θ=θ''}
=2θ+θ'. ----→(a)
 
We know the equation of kinematics,
 
Put all the values in this equation 
=0 x 10 +(1/2 x 4 x 10 x 10)
θ=200 radians
Now, 

=(0+4 x 10)10
θ'=400 radians
Hence the rotation of total angle , by equation (a)
=2θ+θ' = (2 x 200 +400)
=800 rad .